# A Triple-Inertial Accelerated Alternating Optimization Method for Deep Learning Training

### v1

@@ Abstract
The stochastic gradient descent (SGD) algorithm has achieved remarkable success in training deep learning models. However, it has several limitations, including susceptibility to vanishing gradients, sensitivity to input data, and a lack of robust theoretical guarantees. In recent years, alternating minimization (AM) methods have emerged as a promising alternative for model training by employing gradient-free approaches to iteratively update model parameters. Despite their potential, these methods often exhibit slow convergence rates. To address this challenge, we propose a novel Triple-Inertial Accelerated Alternating Minimization (TIAM) framework for neural network training. The TIAM approach incorporates a triple-inertial acceleration strategy with a specialized approximation method, facilitating targeted acceleration of different terms in each sub-problem optimization. This integration improves the efficiency of convergence, achieving superior performance with fewer iterations. Additionally, we provide a convergence analysis of the TIAM algorithm, including its global convergence properties and convergence rate. Extensive experiments validate the effectiveness of the TIAM method, showing significant improvements in generalization capability and computational efficiency compared to existing approaches, particularly when applied to the rectified linear unit (ReLU) and its variants.
@@ 1 Introduction
To overcome the challenges faced by gradient descent algorithms in neural network training, the alternating minimization (AM) method [5] has emerged as a viable alternative for parameter optimization. The AM method decomposes complex global optimization problems into multiple simple sub-problems, reducing overall complexity by iteratively optimizing the variable of corresponding sub-problems in an alternating fashion. This approach provides a flexible and practical optimization strategy, especially in scenarios involving highly structured or large-scale models, such as compressed sensing (CS) [6] and image processing [7]. In this way, model training approaches utilizing AM circumvent the reliance on gradient-based backpropagation, effectively mitigating the vanishing gradient problem. Furthermore, these methods demonstrate better robustness in handling non-smooth regularization terms. Recognizing these advantages, researchers have increasingly focused on the theoretical underpinnings of such methods. For instance, Miguel et al. [8] proposed a method for solving constrained optimization problems, applicable to penalty-based approaches by alternating optimization over the parameters and auxiliary coordinates. Based on the block coordinate descent (BCD) method, Lau et al. [9] proposed a novel algorithm and proved its global convergence, built upon the powerful framework of the Kurdyka-Lojasiewicz (KL) property. Zhang et al. [10] prove that the BCD algorithm globally converges to a stationary point with an R-linear convergence rate of order one. For stochastic or mini-batch strategies, Choromanska et al. [11] proposed a novel online (stochastic/mini-batch) AM approach for training deep neural networks, with the first theoretical convergence guarantees for AM in stochastic settings. Furthermore, Tang et al. [12] proposed a distributed paralleled algorithm regarding recurrent neural networks based on the alternating direction method of multipliers (ADMM) framework and proved relevant convergence.
@@
    Performance Evaluation and Generalization: Extensive experiments were conducted on four benchmark datasets, where the proposed method outperforms state-of-the-art algorithms in terms of test accuracy, computational efficiency, and robustness. Furthermore, most previous studies have focused on the rectified linear unit (ReLU) activation function, with limited experimental exploration of its variants. The proposed algorithm demonstrates superior generalization capability and robustness for ReLU activation functions and their variants, which have been rarely studied experimentally in prior algorithms.
@@ 3.2 Triple-Inertial Acceleration for Alternating Optimization Method
    To the best of our knowledge, most existing AM methods for solving MLP problems primarily utilize the ReLU activation function in experiments, with limited exploration of its variants. In this study, we perform a comprehensive comparison of various ReLU variants to demonstrate the superior performance of our algorithm on ReLU-type activation functions.
@@ 3.3.1 Update $W_{l}$
$P_{l}^{k+1}\left(W_{l};\theta_{l}^{k+1}\right)$ $=\phi\left(\overline{a}_{l-1}^{k+1},\tilde{W}_{l}^{k+1},\overline{z}_{l}^{k},\overline{b}_{l}^{k}\right)+\left(\nabla_{{W}_{l}}\phi\left(\overline{a}_{l-1}^{k+1},\hat{W}_{l}^{k+1},\overline{z}_{l}^{k},\overline{b}_{l}^{k}\right)\right)^{T}$ (5)
$\left(W_{l}-\tilde{W}_{l}^{k+1}\right)+\frac{\theta_{l}^{k+1}}{2}\|W_{l}-\tilde{W}_{l}^{k+1}\|^{2}.$
@@ 3.3.2 Update $b_{l}$
$U_{l}^{k+1}\left(b_{l};\xi_{l}^{k+1}\right)$ $=\phi\left(\overline{a}_{l-1}^{k+1},\overline{W}_{l}^{k+1},\overline{z}_{l}^{k},\tilde{b}_{l}^{k+1}\right)+\left(\nabla_{{b}_{l}}\phi\left(\overline{a}_{l-1}^{k+1},\overline{W}_{l}^{k+1},\overline{z}_{l}^{k},\hat{b}_{l}^{k+1}\right)\right)^{T}$ (9)
$\left(b_{l}-\tilde{b}_{l}^{k+1}\right)+\frac{\xi_{l}^{k+1}}{2}\|b_{l}-\tilde{b}_{l}^{k+1}\|^{2},$
@@
$b_{l}^{k+1}\leftarrow\tilde{b}_{l}^{k+1}-\nabla_{{b}_{l}}\phi\left(\overline{a}_{l-1}^{k+1},\overline{W}_{l}^{k+1},\overline{z}_{l}^{k},\hat{b}_{l}^{k+1}\right)/\xi_{l}^{k+1}$ (12)
@@ 3.3.3 Update $z_{l}$
$V_{l}^{k+1}(z_{l})=$ $\phi\left(\overline{a}_{l-1}^{k+1},\overline{W}_{l}^{k+1},\tilde{z}_{l}^{k+1},\overline{b}_{l}^{k+1}\right)+\left(\nabla_{{z}_{l}}\phi\left(\overline{a}_{l-1}^{k+1},\overline{W}_{l}^{k+1},\hat{z}_{l}^{k+1},\overline{b}_{l}^{k+1}\right)\right)^{T}\left(z_{l}-\tilde{z}_{l}^{k+1}\right)$ (15)
$+\frac{\rho}{2}\|z_{l}-\tilde{z}_{l}^{k+1}\|^{2},$
@@ 3.3.4 Update $a_{l}$
$Q_{l}^{k+1}\left(a_{l};\tau_{l}^{k+1}\right)=$ $\phi\left(\tilde{a}_{l}^{k+1},\overline{W}_{l+1}^{k},\overline{z}_{l+1}^{k},\overline{b}_{l+1}^{k}\right)+\left(\nabla_{{a}_{l}}\phi\left(\hat{a}_{l}^{k+1},\overline{W}_{l+1}^{k},\overline{z}_{l+1}^{k},\overline{b}_{l+1}^{k}\right)\right)^{T}\left(a_{l}-\tilde{a}_{l}^{k+1}\right)$ (20)
$+\frac{\tau_{l}^{k+1}}{2}\|a_{l}-\tilde{a}_{l}^{k+1}\|^{2},$
@@ 4.1 Convergence Properties
(Subgradient Bound) In Algorithm 2, it exist that:
$C_{2}=\max\left(M,M+\theta_{1}^{k+1},M+\theta_{2}^{k+1},\cdots,M+\theta_{L}^{k+1}\right)$, and $g_{2}^{k+1}\in\partial_{\mathbf{W}^{k+1}}F$ such that for any $k\in\mathbb{N}$,
$\|g_{2}^{k+1}\|\leqslant C_{2}\Big{(}\|\mathbf{W}^{k+1}-\mathbf{W}^{k}\|+\|\mathbf{z}^{k+1}-\mathbf{z}^{k}\|+\|\mathbf{a}^{k+1}-\mathbf{a}^{k}\|+\|\mathbf{b}^{k+1}-\mathbf{b}^{k}\|+\|\mathbf{W}^{k}-\mathbf{W}^{k-1}\|+\|\mathbf{b}^{k}-\mathbf{b}^{k-1}\|+\|\mathbf{z}^{k}-\mathbf{z}^{k-1}\|\Big{)}.$
@@ 5 Experiments
Following the experimental settings of the existing algorithm, we constructed a fully connected neural network as our base model. The network consists of three layers, each containing 100 hidden units, with the ReLU activation function [35] applied to each neuron. The selected loss function is cross-entropy with softmax, and the total number of training epochs is set to 200. To mitigate the effects of random initialization of weights and biases, we conducted experiments using 10 different random seeds ranging from 0 to 9. The final evaluation results are presented as the mean and standard deviation of the corresponding epochs, and all the comparison algorithms follow the same settings. The experiments were implemented in Python 3.8 with Torch 1.9 [36] and conducted on a Windows 10 computer equipped with an Intel(R) Core(TM) i7-12700 2.10 GHz CPU and 16GB of memory.
@@ 5.2 Datasets
Citeseer [37]: The Citeseer dataset consists of 3312 scientific papers sourced from the Tagged.com social network. It is classified into 6 categories, with 4732 citation links and a vocabulary of 3703 unique words (i.e. feature).
@@ 6 The Performance of Variants of the ReLU Activation Function
Figures 5, 6, and 7 illustrate the performance of the TIAM algorithm using different ReLU variants. The results show that the ReLU variants achieve better convergence performance than the standard ReLU activation function on the Cora dataset. However, their test performance on the Citeseer dataset is slightly lower than that of the standard ReLU, yet it still surpasses existing AM-based algorithms. Similarly, altering the activation function has little impact on test performance for the Amazon Photo dataset, as all variants outperform other algorithms within 40 iterations and exhibit minimal fluctuations. For the Coauthor CS dataset, although all activation functions converge slightly more slowly in the early stages, all ReLU variants achieve superior test performance later and surpass other algorithms within 75 iterations. Overall, the three ReLU variants perform well across datasets. These findings suggest that the proposed algorithm exhibits superior generalization capability and robustness across different activation functions.
@@ 7 Conclusion
In this work, we proposed a novel AM method with a triple-inertial acceleration framework to solve the MLP optimization problem. Our algorithm combines a specialized approximation approach with a triple-inertial acceleration strategy to enable block-wise accelerated optimization of neural network parameters. Furthermore, we conducted a rigorous convergence analysis and proved that the proposed algorithm achieves a linear convergence rate. Experimental results confirm the superior performance and efficiency of our method. In future work, we plan to incorporate stochastic algorithms into the current framework by adopting a stochastic mini-batch approach, which aims further to enhance the efficiency of AM-based neural network optimization algorithms.
Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgment
This research is supported by the National Natural Science Foundation of China (NSFC) grants 92473208, 12401415, the Key Program of National Natural Science of China 12331011, the 111 Project (No. D23017), the Natural Science Foundation of Hunan Province (No. 2025JJ60009), and the Postgraduate Scientific Research Innovation Project of Hunan Province (No. LXBZZ2024114).
@@ B. Proof of Lemma
Therefore, there exists $g_{2,l}^{k+1}\in\partial_{W_{l}^{k+1}}F$ such that
$\|g_{2,l}^{k+1}\|$ $\leqslant\left(M+\theta_{l}^{k+1}\right)\|W_{l}^{k+1}-W_{l}^{k}\|+\left(M+\theta_{l}^{k+1}\right)\|W_{l}^{k}-W_{l}^{k-1}\|+M\|b_{l}^{k+1}-b_{l}^{k}\|$
@@
So there exists $g_{2}^{k+1}=g_{2,1}^{k+1}\times g_{2,2}^{k+1}\times\cdots\times g_{2,L}^{k+1}\in\partial_{\mathbf{W}^{k+1}}F$, and the coefficient of the corresponding subgradient is $C_{2}=\max\left(M,M+\theta_{1}^{k+1},M+\theta_{2}^{k+1},\cdots,M+\theta_{L}^{k+1}\right)$ such that
$\|g_{2}^{k+1}\|$ $\leqslant C_{2}\Big{(}\|\mathbf{W}^{k+1}-\mathbf{W}^{k}\|+\|\mathbf{z}^{k+1}-\mathbf{z}^{k}\|+\|\mathbf{a}^{k+1}-\mathbf{a}^{k}\|+\|\mathbf{b}^{k+1}-\mathbf{b}^{k}\|+\|\mathbf{W}^{k}-\mathbf{W}^{k-1}\|+\|\mathbf{b}^{k}-\mathbf{b}^{k-1}\|+\|\mathbf{z}^{k}-\mathbf{z}^{k-1}\|\Big{)}.$ (31)
@@
$\|g_{2,l}^{k+1}\|$ $\leqslant\left(M+\theta_{l}^{k+1}\right)\|W_{l}^{k+1}-W_{l}^{k}\|+M\|z_{l}^{k+1}-z_{l}^{k}\|+M\|b_{l}^{k+1}-b_{l}^{k}\|.$ (32)
By combining (31) and (32), we obtain $C_{2}=\max\left(M,M+\theta_{1}^{k+1},M+\theta_{2}^{k+1},\cdots,M+\theta_{L}^{k+1}\right)$ and $g_{2}^{k+1}=g_{2,1}^{k+1}\times g_{2,2}^{k+1}\times\cdots\times g_{2,L}^{k+1}\in\partial_{\mathbf{W}^{k+1}}F$ such that
$\|g_{2}^{k+1}\|$ $\leqslant C_{2}\Big{(}\|\mathbf{W}^{k+1}-\mathbf{W}^{k}\|+\|\mathbf{z}^{k+1}-\mathbf{z}^{k}\|+\|\mathbf{a}^{k+1}-\mathbf{a}^{k}\|+\|\mathbf{b}^{k+1}-\mathbf{b}^{k}\|+\|\mathbf{W}^{k}-\mathbf{W}^{k-1}\|+\|\mathbf{b}^{k}-\mathbf{b}^{k-1}\|+\|\mathbf{z}^{k}-\mathbf{z}^{k-1}\|\Big{)}.$ (33)

### v2

@@ Abstract
The stochastic gradient descent (SGD) algorithm has achieved remarkable success in training deep learning models. However, it has several limitations, including susceptibility to vanishing gradients, sensitivity to input data, and a lack of robust theoretical guarantees. In recent years, alternating minimization (AM) methods have emerged as a promising alternative for model training by employing gradient-free approaches to iteratively update model parameters. Despite their potential, these methods often exhibit slow convergence rates. To address this challenge, we propose a novel Triple-Inertial Accelerated Alternating Minimization (TIAM) framework for neural network training. The TIAM approach incorporates a triple-inertial acceleration strategy with a specialized approximation method, facilitating targeted acceleration of different terms in each sub-problem optimization. This integration improves the efficiency of convergence, achieving better performance with fewer iterations. Additionally, we provide a convergence analysis of the TIAM algorithm, including its global convergence properties and convergence rate. Extensive experiments validate the effectiveness of the TIAM method, demonstrating improvements in generalization capability and computational efficiency compared to existing approaches, particularly when applied to the rectified linear unit (ReLU) and its variants.
@@ 1 Introduction
To overcome the challenges faced by gradient descent algorithms in neural network training, the alternating minimization (AM) method [5] has emerged as a viable alternative for parameter optimization. The AM method decomposes complex global optimization problems into multiple simple sub-problems, reducing overall complexity by iteratively optimizing the variable of corresponding sub-problems in an alternating fashion. This approach provides a flexible and practical optimization strategy, especially in scenarios involving highly structured or large-scale models, such as compressed sensing (CS) [6] and image processing [7, 8]. In this way, model training approaches utilizing AM circumvent the reliance on gradient-based backpropagation, effectively mitigating the vanishing gradient problem. Furthermore, these methods demonstrate better robustness in handling non-smooth regularization terms. Recognizing these advantages, researchers have increasingly focused on the theoretical underpinnings of such methods. For instance, Miguel et al. [9] proposed a method for solving constrained optimization problems, applicable to penalty-based approaches by alternating optimization over the parameters and auxiliary coordinates. Based on the block coordinate descent (BCD) method, Lau et al. [10] proposed a novel algorithm and proved its global convergence, built upon the powerful framework of the Kurdyka-Lojasiewicz (KL) property. Zhang et al. [11] prove that the BCD algorithm globally converges to a stationary point with an R-linear convergence rate of order one. For stochastic or mini-batch strategies, Choromanska et al. [12] proposed a novel online (stochastic/mini-batch) AM approach for training deep neural networks, with the first theoretical convergence guarantees for AM in stochastic settings. Furthermore, Tang et al. [13] proposed a distributed paralleled algorithm regarding recurrent neural networks based on the alternating direction method of multipliers (ADMM) framework and proved relevant convergence.
@@
    Performance Evaluation and Generalization: Extensive experiments were conducted on four benchmark datasets, where the proposed method outperforms state-of-the-art algorithms in terms of test accuracy, computational efficiency, and robustness. Furthermore, most previous studies have focused on the rectified linear unit (ReLU) activation function, with limited experimental exploration of its variants. The proposed algorithm demonstrates better generalization capability and robustness for ReLU activation functions and their variants, which have been rarely studied experimentally in prior algorithms.
@@ 3.2 Triple-Inertial Acceleration for Alternating Optimization Method
    To the best of our knowledge, most existing AM methods for solving MLP problems primarily utilize the ReLU activation function in experiments, with limited exploration of its variants. In this study, we perform a comprehensive comparison of various ReLU variants to demonstrate the better performance of our algorithm on ReLU-type activation functions.
@@ 3.3.1 Update $W_{l}$
$P_{l}^{k+1}\left(W_{l};\theta_{l}^{k+1}\right)$ $=\phi\left(\overline{a}_{l-1}^{k+1},\tilde{W}_{l}^{k+1},\overline{z}_{l}^{k},\overline{b}_{l}^{k}\right)+\left(\nabla_{{W}_{l}}\phi\left(\overline{a}_{l-1}^{k+1},\hat{W}_{l}^{k+1},\overline{z}_{l}^{k},\overline{b}_{l}^{k}\right)\right)^{T}\left(W_{l}-\tilde{W}_{l}^{k+1}\right)+\frac{\theta_{l}^{k+1}}{2}\|W_{l}-\tilde{W}_{l}^{k+1}\|^{2}.$ (5)
@@ 3.3.2 Update $b_{l}$
$U_{l}^{k+1}\left(b_{l};\xi_{l}^{k+1}\right)$ $=\phi\left(\overline{a}_{l-1}^{k+1},\overline{W}_{l}^{k+1},\overline{z}_{l}^{k},\tilde{b}_{l}^{k+1}\right)+\left(\nabla_{{b}_{l}}\phi\left(\overline{a}_{l-1}^{k+1},\overline{W}_{l}^{k+1},\overline{z}_{l}^{k},\hat{b}_{l}^{k+1}\right)\right)^{T}\left(b_{l}-\tilde{b}_{l}^{k+1}\right)+\frac{\xi_{l}^{k+1}}{2}\|b_{l}-\tilde{b}_{l}^{k+1}\|^{2},$ (9)
@@
$b_{l}^{k+1}\leftarrow\tilde{b}_{l}^{k+1}-\nabla_{{b}_{l}}\phi\left(\overline{a}_{l-1}^{k+1},\overline{W}_{l}^{k+1},\overline{z}_{l}^{k},\hat{b}_{l}^{k+1}\right)/\xi_{l}^{k+1}.$ (12)
@@ 3.3.3 Update $z_{l}$
$V_{l}^{k+1}(z_{l})=$ $\phi\left(\overline{a}_{l-1}^{k+1},\overline{W}_{l}^{k+1},\tilde{z}_{l}^{k+1},\overline{b}_{l}^{k+1}\right)+\left(\nabla_{{z}_{l}}\phi\left(\overline{a}_{l-1}^{k+1},\overline{W}_{l}^{k+1},\hat{z}_{l}^{k+1},\overline{b}_{l}^{k+1}\right)\right)^{T}\left(z_{l}-\tilde{z}_{l}^{k+1}\right)+\frac{\rho}{2}\|z_{l}-\tilde{z}_{l}^{k+1}\|^{2},$ (15)
@@ 3.3.4 Update $a_{l}$
$Q_{l}^{k+1}\left(a_{l};\tau_{l}^{k+1}\right)=$ $\phi\left(\tilde{a}_{l}^{k+1},\overline{W}_{l+1}^{k},\overline{z}_{l+1}^{k},\overline{b}_{l+1}^{k}\right)+\left(\nabla_{{a}_{l}}\phi\left(\hat{a}_{l}^{k+1},\overline{W}_{l+1}^{k},\overline{z}_{l+1}^{k},\overline{b}_{l+1}^{k}\right)\right)^{T}\left(a_{l}-\tilde{a}_{l}^{k+1}\right)+\frac{\tau_{l}^{k+1}}{2}\|a_{l}-\tilde{a}_{l}^{k+1}\|^{2},$ (20)
@@ 4.1 Convergence Properties
(Subgradient Bound) In Algorithm 2, it exist that: $C_{1}=\max\left(M,M+\theta_{1}^{k+1},M+\theta_{2}^{k+1},\cdots,M+\theta_{L}^{k+1}\right)$, and $g_{1}^{k+1}\in\partial_{\mathbf{W}^{k+1}}F$ such that for any $k\in\mathbb{N}$,
$\|g_{1}^{k+1}\|\leqslant C_{1}\Big{(}\|\mathbf{W}^{k+1}-\mathbf{W}^{k}\|+\|\mathbf{z}^{k+1}-\mathbf{z}^{k}\|+\|\mathbf{a}^{k+1}-\mathbf{a}^{k}\|+\|\mathbf{b}^{k+1}-\mathbf{b}^{k}\|+\|\mathbf{W}^{k}-\mathbf{W}^{k-1}\|+\|\mathbf{b}^{k}-\mathbf{b}^{k-1}\|+\|\mathbf{z}^{k}-\mathbf{z}^{k-1}\|\Big{)}.$
@@ 5 Experiments
Following the experimental settings of the existing algorithm, we constructed a fully connected neural network as our base model. The network consists of three layers, each containing 100 hidden units, with the ReLU activation function [36] applied to each neuron. The selected loss function is cross-entropy with softmax, and the total number of training epochs is set to 200. To mitigate the effects of random initialization of weights and biases, we conducted experiments using 10 different random seeds ranging from 0 to 9. The final evaluation results are presented as the mean and standard deviation of the corresponding epochs, and all the comparison algorithms follow the same settings. The experiments were implemented in Python 3.8 with Torch 1.9 [37] and conducted on a Windows 10 computer equipped with an Intel(R) Core(TM) i7-12700 2.10 GHz CPU and 16GB of memory. The implementation will be publicly accessible at https://github.com/yancc103/TIAM once this manuscript is accepted.
@@ 5.2 Datasets
Citeseer [38]: The Citeseer dataset consists of 3327 scientific papers sourced from the Tagged.com social network. It is classified into 6 categories, with 4732 citation links and a vocabulary of 3703 unique words (i.e. feature).
@@ 6 The Performance of Variants of the ReLU Activation Function
Figures 5, 6, and 7 illustrate the performance of the TIAM algorithm using different ReLU variants. The results show that the ReLU variants achieve better convergence performance than the standard ReLU activation function on the Cora dataset. However, their test performance on the Citeseer dataset is slightly lower than that of the standard ReLU, yet it still surpasses existing AM-based algorithms. Similarly, altering the activation function has little impact on test performance for the Amazon Photo dataset, as all variants outperform other algorithms within 40 iterations and exhibit minimal fluctuations. For the Coauthor CS dataset, although all activation functions converge slightly more slowly in the early stages, all ReLU variants achieve better test performance later and surpass other algorithms within 75 iterations. Overall, the three ReLU variants perform well across datasets. These findings suggest that the proposed algorithm exhibits better generalization capability and robustness across different activation functions.
@@ 7 Conclusion
In this work, we proposed a novel AM method with a triple-inertial acceleration framework to solve the MLP optimization problem. Our algorithm combines a specialized approximation approach with a triple-inertial acceleration strategy to enable block-wise accelerated optimization of neural network parameters. Furthermore, we conducted a rigorous convergence analysis and proved that the proposed algorithm achieves a linear convergence rate. Experimental results confirm the better performance and efficiency of our method. In future work, we plan to incorporate stochastic algorithms into the current framework by adopting a stochastic mini-batch approach, which aims further to enhance the efficiency of AM-based neural network optimization algorithms.
Declarations
Funding: This research is supported by the National Natural Science Foundation of China (NSFC) grants 92473208, 12401415, the Key Program of National Natural Science of China 12331011, the 111 Project (No. D23017), the Natural Science Foundation of Hunan Province (No. 2025JJ60009), and the Postgraduate Scientific Research Innovation Project of Hunan Province (No. LXBZZ2024114).
Data Availability: Enquiries about data/code availability should be directed to the authors.
Competing interests: The authors have no competing interests to declare that are relevant to the content of this paper.
@@ B. Proof of Lemma
Therefore, there exists $g_{1,l}^{k+1}\in\partial_{W_{l}^{k+1}}F$ such that
$\|g_{1,l}^{k+1}\|$ $\leqslant\left(M+\theta_{l}^{k+1}\right)\|W_{l}^{k+1}-W_{l}^{k}\|+\left(M+\theta_{l}^{k+1}\right)\|W_{l}^{k}-W_{l}^{k-1}\|+M\|b_{l}^{k+1}-b_{l}^{k}\|$
@@
So there exists $g_{1}^{k+1}=g_{1,1}^{k+1}\times g_{1,2}^{k+1}\times\cdots\times g_{1,L}^{k+1}\in\partial_{\mathbf{W}^{k+1}}F$, and the coefficient of the corresponding subgradient is $C_{1}=\max\left(M,M+\theta_{1}^{k+1},M+\theta_{2}^{k+1},\cdots,M+\theta_{L}^{k+1}\right)$ such that
$\|g_{1}^{k+1}\|$ $\leqslant C_{1}\Big{(}\|\mathbf{W}^{k+1}-\mathbf{W}^{k}\|+\|\mathbf{z}^{k+1}-\mathbf{z}^{k}\|+\|\mathbf{a}^{k+1}-\mathbf{a}^{k}\|+\|\mathbf{b}^{k+1}-\mathbf{b}^{k}\|+\|\mathbf{W}^{k}-\mathbf{W}^{k-1}\|+\|\mathbf{b}^{k}-\mathbf{b}^{k-1}\|+\|\mathbf{z}^{k}-\mathbf{z}^{k-1}\|\Big{)}.$ (31)
@@
$\|g_{1,l}^{k+1}\|$ $\leqslant\left(M+\theta_{l}^{k+1}\right)\|W_{l}^{k+1}-W_{l}^{k}\|+M\|z_{l}^{k+1}-z_{l}^{k}\|+M\|b_{l}^{k+1}-b_{l}^{k}\|.$ (32)
By combining (31) and (32), we obtain $C_{1}=\max\left(M,M+\theta_{1}^{k+1},M+\theta_{2}^{k+1},\cdots,M+\theta_{L}^{k+1}\right)$ and $g_{1}^{k+1}=g_{1,1}^{k+1}\times g_{1,2}^{k+1}\times\cdots\times g_{1,L}^{k+1}\in\partial_{\mathbf{W}^{k+1}}F$ such that
$\|g_{1}^{k+1}\|$ $\leqslant C_{1}\Big{(}\|\mathbf{W}^{k+1}-\mathbf{W}^{k}\|+\|\mathbf{z}^{k+1}-\mathbf{z}^{k}\|+\|\mathbf{a}^{k+1}-\mathbf{a}^{k}\|+\|\mathbf{b}^{k+1}-\mathbf{b}^{k}\|+\|\mathbf{W}^{k}-\mathbf{W}^{k-1}\|+\|\mathbf{b}^{k}-\mathbf{b}^{k-1}\|+\|\mathbf{z}^{k}-\mathbf{z}^{k-1}\|\Big{)}.$ (33)